OpenStudy (anonymous):

The sum of the measures of angle X and angle Y is 90. If the measure of angle X is 30 less than twice the measure of angle Y, what is the measure of angle X? 20° 35° 50° 65°

OpenStudy (mathstudent55):

x = 2y - 30 y + 2y - 30 = 90 3y = 120 y = 40 x = 2(40) - 30 x = 50

OpenStudy (anonymous):

where did you get the extra y?

OpenStudy (mathstudent55):

The problem states the sum of the measures of x and y is 90. One angle is y. The other angle is x which is 2y - 30. When you add angle y to angle 2y - 30 you get (y) + (2y -30) = 90 ------- --------- ---- this is y this is x sum

OpenStudy (anonymous):

oh ok :) i see it now Thanks soo much

OpenStudy (mathstudent55):

You're welcome.